Which of the following quadratic equations has solutions x=6a and x=-3b?
A. x^2 - 18ab=0
B. x^2 - x(3b-6a)-18ab=0
C. x^2 - x(3b+6a) +18ab= 0
D. x^2 + x(3b-6a)- 18ab=0
E. x^2 + x(3b+6a) + 18ab=0

Found 2 solutions by Alan3354, drk:
Answer by Alan3354(69443)

Which of the following quadratic equations has solutions x=6a and x=-3b?
A. x^2 - 18ab=0
B. x^2 - x(3b-6a)-18ab=0
C. x^2 - x(3b+6a) +18ab= 0
D. x^2 + x(3b-6a)- 18ab=0
E. x^2 + x(3b+6a) + 18ab=0
---------------------
Multiply it and see.
(x-6a)*(x+3b) = x^2 + x(6a-3b) - 18ab
it's B

Answer by drk(1908)

From our solutions, we will work backward.
x=6a
x=-3b
(x-6a)(x+3b) = 0
x^2 + x(3b-6a) - 18ab.
This is answer [D]